OpenStudy (abdela25):

write an equation in point-slope form for the line that has the given slope 'm' and that passes through the given point m=-2; (0,1)

OpenStudy (anonymous):

y=mx+b now start substituting

OpenStudy (anonymous):

y=(-2)*x+b y=(-2)*0+b

OpenStudy (anonymous):

(-2)(0)+b=1 b=1

OpenStudy (anonymous):

y=-2x+1

OpenStudy (abdela25):

thank you very much such a nice and easy explanation medal was awarded @carolyn0727

OpenStudy (anonymous):

Thank you!